In 1990 an individual earned $24,299 and the CPI had a value of 142. The CPI in 2018 had a value of 226. Calculate the value of the individual's income measured in 2018 dollars. Round your answers to two decimal places
Answer #1

Inflation between the years 1990 and 2018

= (CPI of 2018 - CPI of 1990)/CPI of 1990 * 100

= (226 - 142)/142 * 100

= 84/142 *100

= 0.59154 * 100

= 59.154%

Income in 1990 = $24,299 + 59.154% * 24,299

= 24,299(1.59154)

= $38,673.06

Another way to solve it

= CPI of 2018/CPI of 1990 * Income in 1990

= 226/142 * 24,299

= $38,673.06